# Dynamical stability of various convex graphical translators

### v1

@@ 5 Dynamical stability of graphical translators in $\mathbf{R}^{3}$
By using theorem 1.5 in [undefab] (or theorem 2.4 in [undefu]), we see that for each fixed $|x_{2}|<b$,
@@
We briefly explain the proof of theorem 5.1. The idea is again to use general strategy 1. To do so, we must ensure that the forward limit flow is convex, and we can apply Hamilton’s Harnack inequality. However, because our solutions are noncompact, and the initial surface is not necessarily $C^{1}$, it is unclear if one indeed has both convexity, and validity of Harnack inequality. To overcome this, we follow the ideas in [undefl] and construct the solution as a local smooth limit of convex, compact solutions to mean curvature flow.
@@
The proof is a slight modification of proof of corollary 5.2 in [undefl]. By possibly translating $u_{0}$ vertically, we may assume that
@@
which is a contradiction for large tst-s. Thus $\partial_{t}u^{\infty}\leq 1$. The reverse inequality is obtained similarly. Hence we see that $\partial_{t}u^{\infty}=1$, i.e $M^{\infty}_{0}$ is a graphical translator over $\Omega^{2}_{b}$. (157) implies that $M^{\infty}_{0}$ is not contained in $\Omega^{2}_{b^{\prime}}\times\mathbf{R}$ for $b^{\prime}<b$. Then, by the classification result in [undefu], $M^{\infty}_{0}$ is either (tilted) grim reaper plane, or a $\Delta$-wing defined over $\Omega^{2}_{b}$. It is clear that one type of translator cannot lie in between two copies of another, hence $M^{\infty}=\overline{M}$ up to translation.
@@
The bound (169) is derived as follows. Assume the bound is not true. By reflection in $x_{1}$ variable, we may without loss of generality assume that $c_{1}\geq 0$. We first consider when $c_{0}\geq 0$. Then by theorem 1.5 in [undefab] and (165), we have
@@ 6 Dynamical stability of asymptotically cylindrical translators
In this section, we prove dynamical stability of asymptotically cylindrical translators (see definition 1.1). Note that such translators are fully classified by Bamler-Lai [undefd], [undefc], and are either the bowl soliton, or entire graph translators constructed in [undefu] with possible Euclidean factors. The proof sketch is as follows. We first show that under the given conditions, the forward limit must be an asymptotically cylindrical flow. Then by using the classification result in [undefc], we can deduce that it must be a translating flow with the same velocity as the initially given translator. In the case the initial translator is a bowl soliton with possible Euclidean factors, one can use uniqueness result of bowl solitons [undefs] to show that the two flows are the same up to translation. In the case of flying wing solutions, we show that the two flows are the same up to translation by using the parametrization of space of all non-collapsed translators given in theorem 1.6 in [undefc].
@@
By classical regularity theory of Ecker and Huisken [undefn], [undefm], one obtains a longtime solution to graphical mean curvature flow with $u_{0}$ as the initial data by taking the limit of solutions to the following initial-boundary value problem
@@
Therefore by avoidance principle [undefaf], we have
@@
Since we have uniform local $C^{0}$ bound on $u^{j}$ independent of $j$, by interior regularity theory [undefm], we also have uniform local $C^{m}$ bound on $u^{j}$ independent of $j$. Thus, by possibly passing through subsequence, we can take a local smooth limit
@@
To show that $M^{\infty}_{t}$ is an asymptotically $(n,k)$-cylindrical flow, we follow the ideas of White [undefag]. We consider the parabolic blowdown of the limit flow, namely for $\lambda\to 0$, define
@@
First note that since $M_{0}$ is mean convex, so is $M^{\infty}_{t}$, and $N^{\lambda}_{t}$. Moreover, since $N^{\lambda}_{t}$ is an entire graph, the collection of all translations of $N^{\lambda}_{t}$ in $e_{n+1}$ direction foliates $\mathbf{R}^{n+1}$. Therefore, by standard calibration argument (see [undeft], remark 2.6), we see that $N^{\lambda}_{t}$ has polynomial volume growth, i.e
@@
is a convex foliation of $\mathbf{R}^{n+1}$. Estimate (197) implies that $N^{\lambda}_{t}$ is trapped between $\lambda\overline{M}_{\lambda^{-2}t}\pm C_{0}\lambda e_{n+1}$. Since $N^{\lambda}_{t}$ is an entire graph, and $\lambda\overline{M}_{\lambda^{-2}t}\pm C_{0}\lambda e_{n+1}\to M^{n,k}_{t}$ locally smoothly, we can apply standard calibration argument (see [undeft], remark 2.6) to compare
@@
Thus, $\mu_{t}$ is equal to $M^{n,k}_{t}$ with possible multiplicities. By once again following the ideas of White [undefag], we can rule out high multiplicities. Since the argument is straightforward and is very similar to the previous argument, we record it here in detail. Note that $N^{\lambda}_{t}$ is also a mean convex entire graph. Therefore, we can repeat the previous calibration argument with role of $N^{\lambda}_{t}$ and $\lambda\overline{M}_{\lambda^{-2}t}$ reversed. More specifically, note that
@@
for all $r>0$. Thus, by dividing by $\omega_{n}r^{n}$ where $\omega_{n}$ is the volume of an $n$-dimensional unit ball, letting $r\to 0$ and using $x_{0}\in M^{n,k}_{t}$, we see that the $n$-dimensional Hausdorff density of $\mu_{t}$ at $x_{0}\in M^{n,k}_{t}$ is bounded by 1. Since $\mu_{t}$ is an integral $n$-varifold with support $M^{n,k}_{t}$, this implies that $\mu_{t}$ is equal to multiplicity one $M^{n,k}_{t}$. Thus by local regularity theorem [undefae], the convergence is smooth, hence $M^{\infty}_{t}$ is an asymptotically $(n,k)$-cylindrical flow, hence it must be a non-collapsed, convex translator of velocity $e_{n+1}$. In particular, we can write
@@
in $\mathbf{R}^{n-k+2}$. By uniqueness of bowl soliton proved in [undefs], the lower dimensional translator factor must also be a nk+1n-k+1-dimensional bowl soliton, hence we see that
@@
We now consider the remaining case when $\overline{M}$ is a flying wing soliton constructed in [undefu] with possible Euclidean factor . In this case, $M^{\infty}$ cannot be a bowl soliton with a Euclidean factor. Indeed, if $M^{\infty}$ was a bowl soliton with a Euclidean factor, then by reversing the role of $M^{\infty}$ and $\overline{M}$, and viewing $M^{\infty}$ as the barrier, previous argument implies that $\overline{M}$ also has to be a bowl soliton with a Euclidean factor which is not the case. Thus, both $M^{\infty}$ and $\overline{M}$ are flying wing solitons with possible Euclidean factors, in particular we can apply proposition 2.2 to both $M^{\infty}_{t}$ and $\overline{M}_{t}$.

### v2

@@ 5 Dynamical stability of graphical translators in $\mathbf{R}^{3}$
By using theorem 1.5 in [undefad] (or theorem 2.4 in [undefv]), we see that for each fixed $|x_{2}|<b$,
@@
We briefly explain the proof of theorem 5.1. The idea is again to use general strategy 1. To do so, we must ensure that the forward limit flow is convex, and we can apply Hamilton’s Harnack inequality. However, because our solutions are noncompact, and the initial surface is not necessarily $C^{1}$, it is unclear if one indeed has both convexity, and validity of Harnack inequality. To overcome this, we follow the ideas in [undefm] and construct the solution as a local smooth limit of convex, compact solutions to mean curvature flow.
@@
The proof is a slight modification of proof of corollary 5.2 in [undefm]. By possibly translating $u_{0}$ vertically, we may assume that
@@
which is a contradiction for large tst-s. Thus $\partial_{t}u^{\infty}\leq 1$. The reverse inequality is obtained similarly. Hence we see that $\partial_{t}u^{\infty}=1$, i.e $M^{\infty}_{0}$ is a graphical translator over $\Omega^{2}_{b}$. (162) implies that $M^{\infty}_{0}$ is not contained in $\Omega^{2}_{b^{\prime}}\times\mathbf{R}$ for $b^{\prime}<b$. Then, by the classification result in [undefv], $M^{\infty}_{0}$ is either (tilted) grim reaper plane, or a $\Delta$-wing defined over $\Omega^{2}_{b}$. It is clear that one type of translator cannot lie in between two copies of another, hence $M^{\infty}=\overline{M}$ up to translation.
@@
The bound (174) is derived as follows. Assume the bound is not true. By reflection in $x_{1}$ variable, we may without loss of generality assume that $c_{1}\geq 0$. We first consider when $c_{0}\geq 0$. Then by theorem 1.5 in [undefad] and (170), we have
@@ 6 Dynamical stability of asymptotically cylindrical translators
In this section, we prove dynamical stability of asymptotically cylindrical translators (see definition 1.1). Note that such translators are fully classified by Bamler-Lai [undefd], [undefc], and are either the bowl soliton, or entire graph translators constructed in [undefv] with possible Euclidean factors. The proof sketch is as follows. We first show that under the given conditions, the forward limit must be an asymptotically cylindrical flow. Then by using the classification result in [undefc], we can deduce that it must be a translating flow with the same velocity as the initially given translator. In the case the initial translator is a bowl soliton with possible Euclidean factors, one can use uniqueness result of bowl solitons [undeft] to show that the two flows are the same up to translation. In the case of flying wing solutions, we show that the two flows are the same up to translation by using the parametrization of space of all non-collapsed translators given in theorem 1.6 in [undefc].
@@
By classical regularity theory of Ecker and Huisken [undefo], [undefn], one obtains a longtime solution to graphical mean curvature flow with $u_{0}$ as the initial data by taking the limit of solutions to the following initial-boundary value problem
@@
Therefore by avoidance principle [undefah], we have
@@
Since we have uniform local $C^{0}$ bound on $u^{j}$ independent of $j$, by interior regularity theory [undefn], we also have uniform local $C^{m}$ bound on $u^{j}$ independent of $j$. Thus, by possibly passing through subsequence, we can take a local smooth limit
@@
To show that $M^{\infty}_{t}$ is an asymptotically $(n,k)$-cylindrical flow, we follow the ideas of White [undefai]. We consider the parabolic blowdown of the limit flow, namely for $\lambda\to 0$, define
@@
First note that since $M_{0}$ is mean convex, so is $M^{\infty}_{t}$, and $N^{\lambda}_{t}$. Moreover, since $N^{\lambda}_{t}$ is an entire graph, the collection of all translations of $N^{\lambda}_{t}$ in $e_{n+1}$ direction foliates $\mathbf{R}^{n+1}$. Therefore, by standard calibration argument (see [undefu], remark 2.6), we see that $N^{\lambda}_{t}$ has polynomial volume growth, i.e
@@
is a convex foliation of $\mathbf{R}^{n+1}$. Estimate (202) implies that $N^{\lambda}_{t}$ is trapped between $\lambda\overline{M}_{\lambda^{-2}t}\pm C_{0}\lambda e_{n+1}$. Since $N^{\lambda}_{t}$ is an entire graph, and $\lambda\overline{M}_{\lambda^{-2}t}\pm C_{0}\lambda e_{n+1}\to M^{n,k}_{t}$ locally smoothly, we can apply standard calibration argument (see [undefu], remark 2.6) to compare
@@
Thus, $\mu_{t}$ is equal to $M^{n,k}_{t}$ with possible multiplicities. By once again following the ideas of White [undefai], we can rule out high multiplicities. Since the argument is straightforward and is very similar to the previous argument, we record it here in detail. Note that $N^{\lambda}_{t}$ is also a mean convex entire graph. Therefore, we can repeat the previous calibration argument with role of $N^{\lambda}_{t}$ and $\lambda\overline{M}_{\lambda^{-2}t}$ reversed. More specifically, note that
@@
for all $r>0$. Thus, by dividing by $\omega_{n}r^{n}$ where $\omega_{n}$ is the volume of an $n$-dimensional unit ball, letting $r\to 0$ and using $x_{0}\in M^{n,k}_{t}$, we see that the $n$-dimensional Hausdorff density of $\mu_{t}$ at $x_{0}\in M^{n,k}_{t}$ is bounded by 1. Since $\mu_{t}$ is an integral $n$-varifold with support $M^{n,k}_{t}$, this implies that $\mu_{t}$ is equal to multiplicity one $M^{n,k}_{t}$. Thus by local regularity theorem [undefag], the convergence is smooth, hence $M^{\infty}_{t}$ is an asymptotically $(n,k)$-cylindrical flow, hence it must be a non-collapsed, convex translator of velocity $e_{n+1}$. In particular, we can write
@@
in $\mathbf{R}^{n-k+2}$. By uniqueness of bowl soliton proved in [undeft], the lower dimensional translator factor must also be a nk+1n-k+1-dimensional bowl soliton, hence we see that
@@
We now consider the remaining case when $\overline{M}$ is a flying wing soliton constructed in [undefv] with possible Euclidean factor . In this case, $M^{\infty}$ cannot be a bowl soliton with a Euclidean factor. Indeed, if $M^{\infty}$ was a bowl soliton with a Euclidean factor, then by reversing the role of $M^{\infty}$ and $\overline{M}$, and viewing $M^{\infty}$ as the barrier, previous argument implies that $\overline{M}$ also has to be a bowl soliton with a Euclidean factor which is not the case. Thus, both $M^{\infty}$ and $\overline{M}$ are flying wing solitons with possible Euclidean factors, in particular we can apply proposition 2.2 to both $M^{\infty}_{t}$ and $\overline{M}_{t}$.
@@
We now prove corollary 1.12. The proof is a slight modification of proof of theorem 6.1, so we only point out the necessary changes.
Proof of corollary 1.12.
The proof is essentially that of theorem 6.1 with slight modifications when proving that $M^{\infty}_{t}$ is asymptotically $(n,k)$-cylindrical. First, by the entropy bound in (ii) of (16), together with the fact that entropy is non-increasing under mean curvature flow, we see that
$\lambda(M^{\infty}_{t})\leq\lambda(M_{0})<2\lambda(\overline{M})\text{ for all }t\in\mathbf{R}.$ (241)
This implies that one can still consider the tangent flow at infinity
$\mu_{t}=\lim_{\lambda\to 0}N^{\lambda}_{t},$ (242)
with $N^{\lambda}_{t}=\lambda M^{\infty}_{\lambda^{-2}t}$. $\mu_{t}$ is currently just an integral $n$-Brakke flow. This tangent flow is self shrinking flow, i.e. $\mu_{t}=\sqrt{-t}\mu_{-1}$, with $\mu_{-1}$ being the pushforward of a $F$-stationary integral varifold, denoted by V(1)V(-1).
By following the same arguments in proof of theorem 6.1, we can show that
$\textup{supp}(\mu_{t})=M^{n,k}_{t}.$ (243)
Note that in this part of the proof, we did not use mean convexity, hence above conclusion is still valid under our alternative assumptions (16). Then since we know that the support of a $F$-stationary varifold V(1)V(-1) is equal to $M^{n,k}_{-1}$, which is a connected, smooth, embedded hypersurface, by constancy theorem for $F$-stationary integral varifolds (see section 8.4 in [undefac]), we see that
$\mu_{-1}=m\mathcal{H}^{n}|_{M^{n,k}_{-1}}$ (244)
for some positive integer $m$. The entropy bound (241) immediately implies that $m=1$, i.e $\mu_{t}$ has multiplicity one. Local regularity theorem [undefag] implies that the convergence (242) is actually smooth, meaning $M^{\infty}_{t}$ is an asymptotically $(n,k)$-cylindrical flow. Then by following the remaining proof of theorem 1.9, we obtain corollary 1.12. ∎